OpenStudy (emilyjones284):

Please help I'll give a medal (: Ill post it below (:

OpenStudy (emilyjones284):

Solve the logarithmic equation by using properties of logs. Round to the nearest thousandth if necessary. \[\log X + \log 4 = -2\]

OpenStudy (campbell_st):

well start by using the log law \[\log(ab) = \log(a) + \log(b)\] so you are now looking at \[\log(4x) = -2\] next raise each side of the equation as a power of 10. the log law is \[a^{\log_{a}(b)} = b\] so \[10^{\log(4x)}= 10^{-2}\] or \[4x = 10^{-2}\] now just solve for x

OpenStudy (emilyjones284):

Thank you! (:

OpenStudy (emilyjones284):

wait i got .0025 and thats not one of the answers

OpenStudy (campbell_st):

what about 1/400